OpenStudy (anonymous):

The Derivative of y=x ln ^{3}x ?

OpenStudy (anonymous):

I think you've made a typo in your equation, can you have another look at it?

OpenStudy (anonymous):

\[y = x \ln ^{3}x\]

OpenStudy (anonymous):

Oh it was right, ignore me haha. So to do this we're going to need the product rule. If we say f(x)=x and g(x)=ln^{3}x, what is f'(x) and g'(x)?

OpenStudy (anonymous):

f'(x) = 1 and g'(x) =\[3x (\ln x)^{2} \]

OpenStudy (anonymous):

Not quite, very close though. d/dx(ln(x))=1/x so it should be 3ln(x)^2/x

OpenStudy (anonymous):

\[3\ln x ^{\frac{ 2 }{ x}} \] Like this??

OpenStudy (anonymous):

nvm i figured out the answer.

OpenStudy (anonymous):

Great :)